OpenStudy (anonymous):

What is the simplified form of the expression? (B/7)^2 A. B^2/ 14 B. B^2 /44 C. B^2 /7 D. 7B^2

OpenStudy (anonymous):

B^2/49

OpenStudy (lgbasallote):

lol again your book fails you =))

OpenStudy (lgbasallote):

the answer is not in the choices :(

OpenStudy (anonymous):

LOL i'm telling you

OpenStudy (anonymous):

B^2/49 is the answer it is not in coices

OpenStudy (lgbasallote):

\[(\frac{b}{7})^2 \implies \frac{b^2}{7^2} \implies \frac{b^2}{49}\]

OpenStudy (anonymous):

Well Im going to talk to the teacher about this.

OpenStudy (lgbasallote):

you go girl! hehe

OpenStudy (anonymous):

Lol it gets on my nerves

OpenStudy (anonymous):

I'm about to post another question myabe you can help

OpenStudy (anonymous):

maybe

OpenStudy (lgbasallote):

sure